# Optimising quantum tomography via shadow inversion

### v3

@@ Introduction
In the previous section we have studied examples of overcomplete POVMs in the entire HS space, i.e. with $D=d^{2}$. Here we will pursue a relevant example of a POVM spanning only a subspace $\mathcal{V}_{D}$ of lower dimension $D<d^{2}$, in which $E_{k}$ forms an over-complete basis (within $\mathcal{V}_{D}$). Note that the standard inversion procedure Huang et al. (2020); Nguyen et al. (2022) is defined on the full HS space and does not directly apply in this instance.
Consider for instance the POVM constituted by the projectors of the X and Y Paulis $E^{\pm}_{\zeta}=\frac{1}{4}(\mathbbm{1}\pm\sigma_{\zeta})$, where $\zeta=x,\,y$, which spans the ”equator” on the Bloch sphere. With a straightforward application of our procedure Eq. (8), we obtain the following particular solution
@@
The homogeneous term is added by solving $\sum_{\zeta}h_{\zeta}^{+}E_{\zeta}^{+}+h_{\zeta}^{-}E_{\zeta}^{-}=0$, where $\{h_{\zeta}\}_{\zeta=x,y}=\{p,-p\}$ and $p$ is a free parameter . For the generic planar observable (within span of $\{E_{\zeta}^{\pm}\}$)
@@
we find the shadow norm
@@
We can now proceed to two examples of estimation of product observables via local measurements achieving the same scaling of resources as global strategies Huang et al. (2020).
@@
When considering as target observable the planar Pauli,
@@
the optimised shadow norm from Eq. (21) results in $\|A\|^{2}_{\text{opt}}=2$. An immediate result is that the expected number of resources for product of planar Paulis scales as $2^{N}$, which matches the scaling obtained from global Clifford measurements Huang et al. (2020), $\|A\|^{2}_{\text{sh}}=\mbox{Tr}\left[{A^{2}}\right]=2^{N}$. Surprisingly, we are able to match the optimal scaling Huang et al. (2021b) while still relying only on local single-qubit measurements.
@@
In the context of the previous example, it is interesting to compare the performance of complete versus overcomplete POVM. For this purpose, consider the example of a triangle POVM on the plane, defined as the normalised projectors $E_{k}=\frac{2}{3}\outerproduct{\psi_{k}}{\psi_{k}}$ | of states
@@
Since the POVM is complete, there is no free parameter dependence. From here, we can compute the shadow norm as defined in Eq.(12). For a planar Pauli as target observable, as in Eq. (22), the shadow norm is $\|A\|^{2}_{\text{sh}}=3,\,\forall\varphi$, which leads to an expected overhead of $3^{N}$ on the scaling of resources. From the previous example, the same observables can be optimally estimated with an overhead of $2^{N}$. We have therefore found an example in which an overcomplete POVM performs better than a complete measurement scheme in the same subspace.
@@
We now present a configuration in which the optimisation of the inversion map not only improves the sample complexity, but is actually able to achieve efficient estimation, i.e. $\|A\|^{2}_{\text{opt}}=1$.
To show this, we choose as target observables tensor products of projectors that act on a qubit, $A=\bigotimes_{j=1}^{N}\outerproduct{\psi(\phi_{j})}{\psi(\phi_{j})}$ | where each single-qubit state is parameterised as: $|{\psi(\phi_{j})}\rangle=(|{0}\rangle+e^{i\phi_{j}}|{1}\rangle)\sqrt{2}$.
Using the change of coordinates
@@
we express the generic shadow norm from Eq. (21) in terms of the parameter $\phi$:
@@
This is particularly relevant considering that X and Y are non-commuting observables, as it opens the possibility of the simultaneous estimation of up to $4^{N}$ projections simultaneously for complementary observables. Previously this result has been found in Huang et al. (2020) only via global measurement strategies relying on $N-$qubit Clifford unitaries. We have shown instead that this is achieved using exclusively local measurements.
@@ Discussion
In the course of this letter we have observed how post-processing, an often overlooked aspect in estimation, represents a relevant resource to be examined in the accounting of a more efficient tomography. By introducing a generalised framework to obtain families of equivalent parameter-dependent estimators, we have found a clear relation between over-complete POVMs and their ability to reduce the variance in estimation of arbitrary observables. When considering local measurements on large composite quantum systems, even a small gain on each estimator represents an overall improvement in sample complexity, yielding an exponential separation between the naive sample variance and the optimised one as seen in Figure 1(b). In the case of example two, this optimised gain can achieve a variance that performs as well as global classical shadows for given observables, with modest additional post processing that remains efficient. We anticipate that by considering overcomplete POVMs containing joint measurements on multiple qubits, further advantage may be found at increased (but still tractable) optimisation difficulty, either in sample complexity or the set of observables. Finally, interleaving optimised multiqubit measurements may yield sample improvements that approach regular classical shadows within some limit. Such questions are natural candidates for further investigation and the identification of such a hierarchy of optimisation will be the topic of future work.
Note - In preparation of this manuscript, we became aware of two other works Malmi et al. (2024); Fischer et al. (2024) considering a similar problem. The authors consider a similar optimisation of the post-processing stage in estimation tasks, albeit referring to different frameworks.
@@ Acknowledgements
We thank Richard Küng for his insight and excellent discussions on the subject matter. This research was funded in whole, or in part, by the Austrian Science Fund (FWF) [10.55776/F71] (BeyondC). For the purpose of open access, the author(s) has applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission.

### v4

@@ Introduction
In the previous section we have studied examples of overcomplete POVMs in the entire HS space, i.e. with $D=d^{2}$. Here we will pursue a relevant example of a POVM spanning only a subspace $\mathcal{V}_{D}$ of lower dimension $D<d^{2}$, in which $E_{k}$ forms an overcomplete basis (within $\mathcal{V}_{D}$). Note that the standard inversion procedure [1, 27] is defined on the full HS space and does not directly apply in this instance.
Consider, for instance, the POVM constituted by the projectors of the $X$ and $Y$ Paulis $E^{\pm}_{\zeta}=\frac{1}{4}(\mathbbm{1}\pm\sigma_{\zeta})$, where $\zeta=x,\,y$, which spans the “equator” on the Bloch sphere. With a straightforward application of our procedure given by Eq. (8), we obtain the following particular solution
@@
The homogeneous term is added by solving $\sum_{\zeta}h_{\zeta}^{+}E_{\zeta}^{+}+h_{\zeta}^{-}E_{\zeta}^{-}=0$, where $\{h_{\zeta}\}_{\zeta=x,y}=\{p,-p\}$ and $p$ is a free parameter . For the generic planar observable (within the span of $\{E_{\zeta}^{\pm}\}$),
@@
we find the shadow norm,
@@
We can now proceed to two examples of estimation of product observables via local measurements achieving the same scaling of resources as global strategies [1].
@@
When considering, as the target observable, the planar Pauli,
@@
the optimized shadow norm from Eq. (21) results in $\|A\|^{2}_{\text{opt}}=2$. An immediate result is that the expected number of resources for the product of planar Pauli eigenstates scales as $2^{N}$, which matches the scaling obtained from global Clifford measurements [1], $\|A\|^{2}_{\text{sh}}=\mbox{Tr}\left[{A^{2}}\right]=2^{N}$. Surprisingly, we are able to match the optimal scaling [44] while still relying only on local single-qubit measurements.
@@
In the context of the previous example, it is interesting to compare the performance of complete versus overcomplete POVM. For this purpose, consider the example of a triangle POVM on the plane, defined as the normalized projectors $E_{k}=\frac{2}{3}\outerproduct{\psi_{k}}{\psi_{k}}$ | of states
@@
Since the POVM is complete, there is no free parameter dependence. From here, we can compute the shadow norm as defined in Eq. (12). For a planar Pauli as the target observable, as in Eq. (22), the shadow norm is $\|A\|^{2}_{\text{sh}}=3,\,\forall\varphi$, which leads to an expected overhead of $3^{N}$ on the scaling of resources. From the previous example, the same observables can be optimally estimated with an overhead of $2^{N}$. We have, therefore, found an example in which an overcomplete POVM performs better than a complete measurement scheme in the same subspace.
@@
We now present a configuration in which the optimization of the inversion map not only improves the sample complexity, but is actually able to achieve efficient estimation, i.e., $\|A\|^{2}_{\text{opt}}=1$.
To show this, we choose, as target observables, tensor products of projectors that act on a qubit, $A=\bigotimes_{j=1}^{N}\outerproduct{\psi(\phi_{j})}{\psi(\phi_{j})}$ |, where each single-qubit state is parameterized as $|{\psi(\phi_{j})}\rangle=(|{0}\rangle+e^{i\phi_{j}}|{1}\rangle)\sqrt{2}$.
Using the change of coordinates,
@@
we express the generic shadow norm from Eq. (21) in terms of the parameter $\phi$,
@@
This is particularly relevant considering that X and Y are non-commuting observables, as it opens the possibility of the simultaneous estimation of up to $4^{N}$ projections simultaneously for complementary observables. Previously, this result has been found in [1] only via global measurement strategies relying on $N-$qubit Clifford unitaries. We have shown instead that this is achieved using exclusively local measurements.
@@ Discussion
In the course of this paper, we have observed how postprocessing, an often overlooked aspect in estimation, represents a relevant resource to be examined in the accounting of a more efficient tomography. By introducing a generalized framework to obtain families of equivalent parameter-dependent estimators, we have found a clear relation between overcomplete POVMs and their ability to reduce the variance in the estimation of arbitrary observables. When considering local measurements on large composite quantum systems, even a small gain on each estimator represents an overall improvement in sample complexity, yielding an exponential separation between the naive sample variance and the optimized one as seen in Fig. 1(b). In the case of example two, this optimized gain can achieve a variance that performs as well as global classical shadows for given observables, with modest additional post processing that remains efficient. We anticipate that by considering overcomplete POVMs containing joint measurements on multiple qubits, further advantage may be found at increased (but still tractable) optimization difficulty, either in sample complexity or the set of observables. Finally, interleaving optimized multiqubit measurements may yield sample improvements that approach regular classical shadows within some limit. Such questions are natural candidates for further investigation and the identification of such a hierarchy of optimization will be the topic of future work.
Note added - Recently, we became aware of two other works [45, 46] considering a similar problem. The authors consider a similar optimization of the postprocessing stage in estimation tasks, albeit referring to different frameworks.
@@ Acknowledgements
We thank Richard Küng for his insight and excellent discussions on the subject matter. This research was funded in whole or in part by the Austrian Science Fund (FWF)(Grant No.10.55776/F71]) (BeyondC).